OpenStudy (g111):

finding theta of complex number z=(3+j4) /(4-j3)

OpenStudy (g111):

is it theta for each domi and nom

OpenStudy (g111):

the modulus is 1

OpenStudy (nick63689):

I can help you by motivating you with a funny joke,

OpenStudy (nick63689):

OH NO! Xbox and PSN and down, CALL AN AMBULANCE! *Wii U Wii U Wii U Wii U Wii U*

OpenStudy (g111):

Looool \[\theta=\tan^{-1} (-3/4) \] or \[\theta=\tan^{-1} (4/3) \]

OpenStudy (g111):

or tan inv of 1

OpenStudy (irishboy123):

sort out the denominator times it (and numerator) by its conjugate go from there good answer waiting you

OpenStudy (g111):

its j

OpenStudy (g111):

what is theta still cant get that

OpenStudy (irishboy123):

 \(z = x + i y = 0 + i\) \( \tan \theta = \dfrac{y}{x}\) so \(\theta = ?\)